Circular No. 5801

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) Following the remarks on IAUC 5800 concerning the encounter with

Jupiter in July 1994, it should be noted that the 3^{o} difference in p.a. between the comet’s direction of approach and the orientation of the nuclear train should increase immediately before encounter, and the undersigned’s initial estimate is that more than half of the nuclear train could collide with Jupiter—over an interval approaching three days. Preliminary computations by A. Carusi, Rome University, show that surviving nuclei are likely to remain as satellites of Jupiter or be thrown closer to the sun on short-period heliocentric orbits (depending on which side of Jupiter they pass). It must be emphasized that a 1994 collision of the train center with Jupiter is not assured, and in the case of a miss, the 3^{o} difference in p.a. would minimize the chance of collision with any part of the train.

1993 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
May 13 12^{h}07.^{m}45 − 2^{o}04.^{′}0 4.755 5.462 130.^{o}2 8.^{o}1 13.5
23 12 06.15 − 1 54.9 4.880 5.460 120.2  9.2 13.6
June 2 12 05.98 − 1 52.8 5.019 5.458 110.5 10.0 13.7
12 12 06.92 − 1 57.8 5.166 5.455 101.2 10.5 13.8
22 12 08.92 − 2 09.4 5.318 5.453  92.2 10.7 13.8
July 2 12 11.92 − 2 27.5 5.471 5.451  83.5 10.7 13.9
12 12 15.82 − 2 51.4 5.622 5.448  75.0 10.4 13.9
22 12 20.53 − 3 20.5 5.766 5.445  66.7  9.9 14.0
Aug. 1 12 25.96 − 3 54.2 5.901 5.443  58.7  9.2 14.0
11 12 32.03 − 4 31.9 6.025 5.440  50.7  8.3 14.0

The orbit on IAUC 5800 indicates that the comet was already in jovicentric orbit before 1992, but this conclusion is very uncertain. Nevertheless, the following ephemeris may be useful for 1992 searches; the magnitude is, of course, a pure guess.

1992 TT α_{2000} δ_{2000} Δ r ϵ β m_{2}
Feb. 8 10^{h}48.^{m}10 + 6^{o}54.^{′}3 4.572 5.498 157.^{o}8 3.^{o}9 22.4
18 10 43.94 + 7 22.5 4.522 5.496 169.0 2.0 22.2
28 10 39.47 + 7 53.1 4.503 5.493 179.4 0.1 22.0
Mar. 9 10 35.02 + 8 24.0 4.515 5.491 168.4 2.1 22.2
19 10 30.92 + 8 53.1 4.556 5.489 157.4 4.0 22.4
29 10 27.48 + 9 18.7 4.625 5.486 146.5 5.8 22.5
Apr. 8 10 24.92 + 9 39.3 4.719 5.483 135.9 7.3 22.6

1993 May 22 Brian G. Marsden